answersLogoWhite

0

🎒

Probability

The probability of a certain event is a number expressing the likelihood that a specific event will occur, expressed as the ratio of the number of actual occurrences to the number of possible occurrences. In mathematics, it is a measure of how often an event will happen and is the basis of statistics.

14,643 Questions

What is the probability when you take a card from a pack it will be a diamond?

There are 4 suits in a deck of cards; each suit has a probability of being selected of 1/4. So, probability of a diamond is 1/4 or 0.25.

What two ways could a die be tested to see if its fair?

Two ways a die could be tested for fairness are empirical and analytic.

In the empirical method, you throw the die a large number of times and measure the distribution of the results - they should be evenly distributed within whatever level of tolerance (based on number of trials) that you want.

In the analytic method, you measure the dimensions of the die with extremely accurate equipment. This must include checks for physical symmetry and balance. Balance is checked with a balance beam that allows the die to be rotated about various axes to make sure the center of gravity is in the center of the die.

What is the probability of obtaining a number less than 5 when a single fair die is tossed?

How many numbers less than 5 = 4 (1, 2, 3, 4)

How many sides on a fair number cube = 6

Put 4 over 6 = 4/6

Simplify:

4 ÷ 2 = 2

6 ÷ 2 = 3

Making the answer 2/3.

What is the sample size for standard deviation?

There is no such thing. The standard error can be calculated for a sample of any size greater than 1.

What is the probability of having three dauthers in a row?

Each chance of a having a girl is 50% or 1/2. Three different children are separate events and so multiplying the chance all three together gives 1/2 x 1/2 x 1/2= 1/8. 1 in 8 chances will someone have three girls in a row.

What are the odds of drawing a 9 from a deck of cards?

If you don't specify the color, there are four of each card. So out of 52 cards there are four nines. Four out of 52 can be simplified into 1/13, which is about 8%.

What is the probability that Carl and at least one female are given tasks?

The statement of the question doesn't provide sufficient information to calculate that.

What is a real life situation involving subjective probability?

Subjective probability is defined as a person's own judgment of what is or is not going to happen, which does not consist of any kind of mathematical or calculated basis. One situation in which subjective probability would be a factor would be in the stock market. If a person felt that a certain stock would plummet and just had a feeling about it, and sold that stock away, that decision would have been based on subjective probability.

What is the probability that you are dealt a pair or worse?

The answer depends on what game you are playing and so how many cards you are dealt!

What is the probability of spinning a 3 on a four sided spinner?

Assuming that the four-sided spinner is fair and that it is numbered in the traditional way of 1, 2, 3 and 4, the probability of spinning a three is 1/4.

What is the formula to convert normal distribution to the standard normal distribution?

If a variable X, is distributed Normally with mean m and standard deviation s then

Z = (X - m)/s has a standard normal distribution.

What would be the probability that a couple would have 4 sons in succession in a row?

It is not possible to give a proper answer to the question for two main reasons. The first reason is that the probability of boys and girls are not equal. The global probability, at birth is 0.517 for boys and 0.483 for girls. Second, the children's genders are not independent events. Third, the gender ratios change with the parents' (mother's) age.

If you choose to ignore all these facts, then the probability is (1/2)4 = 1/16

What is a probability that a child would be a girl versus a boy?

It would depend on the parents' genes and age. Globally, the probability is approx 0.483

How do you find the interquartile range when given the mean and standard deviation?

In general you cannot. You will need to know more about the distribution of the variable - you cannot assume that the distribution is uniform or Normal.

What is the relationship between the Z-score for Normal Distribution and Continuous Probability?

There is no real relationship. Probabilities for the Normal distribution are extremely difficult to work out. The z-score is a method used to convert any Normal distribution into the Standard Normal distribution so that its probabilities can be looked up in tables easily.

There are infinitely many types of continuous probability distributions and the Normal is just one of them.

What is the probability of chossing a 10 or jack from a standard deck of 52 playing cards?

The probability of choosing a 10 is 4/52; likewise a Jack is 4/52. The probability of 10 or Jack is 4/52 + 4/52 = 8/52 or 2/13.

What is a situation where there are 12 possible outcomes?

William has 4 shirts that he can wear with a choice of 3 different pants. He has 12 different possible outcomes.

Why there are 12 possible outcomes: There are twelve possible outcomes because 4 multiplied by 3 is 12. This is called algorthim. Algorithm is when you multiply numbers to get your outcome.

There are plenty ways to get or find how many outcomes there are. These ways are a tree diagram, chart, sample spaces, ect.

Hope i helped! :)

Are volunteer sampling and self sampling forms of convenience sampling?

No. Convenience sampling involves the active participation of the researcher. It's called convenience sampling because the researcher simply grabs whomever happens to be close to hand. It's not whether people volunteer themselves. Please see the link.

How should a student study during exams?

Time manage so you have slots to revise for each subject! :)

Trending Questions
Once an isolating event occurs a unit's responsibilities inlude? What is the intersection of the assessed probability? When did Odd Eidem die? A number cube was rolled as part of an experiment. The results are shown in the table. Explain how to find the experimental probability of rolling a 6.? In the last 40 days Mr.Neptune's newspaper has been delivered late 6 times what is the probability? What is a theoretical method? How are probability and statistics related? What is the probability of drawing the six clubs from a standard deck of cards? If you select two apples from a bag of 6 red and 8 yellow apples without replacement what is the probability that the second apple you pick will be red? How many 5 numbers combinations can you make from the numbers 1 to 49 and you can use multiple combinations but still use each number only once? What is the probability of getting 2 tails if you tossed 3 coins? What is the probability of drawing a red or black? A Card is drawn from a standard deck of 52 cards What is the probability that the card is either an eight or a face card? What is the joint probability of rolling a four on a six-sided die and correctly quessing a random picked number as being odd or even? A die is rolled what is the probability that the die will show 4 5 or 6? What are the odds of throwing a six on a dice two times? What is the probability of the next card is higher if the card is 7 in 13 cards in an order? What is the probability that a Condorcet winner exists? What is the binomial theorem formula for probability? What is 1standard deviation below 100?